Whitney received $72,300 of taxable income in 2019. All of the income was salary from her employer. What is her income tax liability in each of the following alternative situations? Use Tax Rate Schedule for reference. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

a. She files under the single filing status.

  39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475

b. She files a joint tax return with her spouse. Together their taxable income is $72,300.

$ 19,400 $ 78,950 $1,940 plus 12% of the excess over $19,400

c. She is married but files a separate tax return. Her taxable income is $72,300.

$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475

d. She files as a head of household.

$ 52,850 $ 84,200 $6,065 plus 22% of the excess over $52,850
Answer #1

Using the tax brackets and rates for 2019, tax liability is computed as :

a. She files under the single filing status.

Taxable income = $ 72,300

Tax payable = (72300-39,475) X 22% + 4543 = 7221.5 + 4543 = $ 11764.50

b.She files a joint tax return with her spouse. Together their taxable income is $72,300.

Taxable income = $ 72,300

Tax payable = (72,300 – 19,400) X 12% + 1940 = 6348 + 1940 = $ 8288

c. She is married but files a separate tax return. Her taxable income is $72,300.

Taxable income = $ 72,300

Tax payable = ( 72300 – 39475) X 22% + 4543 = $ 11764.50

d. She files as a head of household

Taxable income = $ 72,300

Tax payable = (72300 – 52850) X 22% + 6065 = 4279 + 6065 = $ 10,344